### v1

@@ 5.3 Validity check of hSIDM approach
Figure 6: Angular distribution $f(t,\theta)\cdot\sin(\theta)$ of beam particles within the deflection test set-up as obtained within the hSIDM scheme at three (very early) times $t$ (solid lines), and for two values of the critical angle $\theta_{\mathrm{c}}=0.3$ (left panel) and $\theta_{\mathrm{c}}=0.5$ (right panel), respectively. The angular distribution is expected to agree with the analytical result Eq. (18) (dashed lines) after some minimal time $t>t_{\rm min}(\theta_{\mathrm{c}})$ for which scatterings by angles $\theta<\theta_{\mathrm{c}}$ are sufficiently frequent such that the effective drag force and momentum diffusion approach applies. The agreement is indeed reached after the expected time $t_{\rm min}(0.3)\simeq 2\cdot 10^{-2}t_{1}$ and $t_{\rm min}(0.5)\simeq 4\cdot 10^{-2}t_{1}$, respectively. We note that hSIDM is thus well applicable on the relevant dynamical time-scale $t\sim t_{\rm dyn}(\equiv t_{1})$.
@@
Figure 7: Minimal time-scale $t_{\rm min}(\theta_{\mathrm{c}})$ after which the hSIDM approach is applicable, extracted from the deflection test set-up. The symbols are derived from a comparison of numerical solutions using the hSIDM scheme with the result obtained when using the numerically expensive explicit sampling technique for all scattering angles (anisotropic rSIDM scheme). We show results for $\theta_{\mathrm{c}}=0.1,0.3,0.5,0.7,0.9$, and for differential cross sections with anisotropy parameter $r=10^{2},10^{3},10^{4}$, respectively. We furthermore consider the case without recoil and with ‘Rutherford’ scattering cross section (upper panel) as well we the case with recoil and with ‘Møller’ cross section (lower panel). Solid lines show the dependence of the analytical prediction Eq. (35) for $t_{\rm min}(\theta_{\mathrm{c}})$ on $\theta_{\mathrm{c}}$ and on $r$.
@@
We note that $t_{\rm min}(\theta_{\mathrm{c}})\ll t_{\rm dyn}(=t_{1})$ is much smaller than the dynamical time-scale for both values of $\theta_{\mathrm{c}}$ shown in Fig. 6. This means the hSIDM approach yields a valid description on the relevant time-scale of order $t_{\rm dyn}$. In the upper panel of Fig. 7 we show $t_{\rm min}(\theta_{\mathrm{c}})$ as obtained from hSIDM simulations with various critical angles $\theta_{\mathrm{c}}=0.1,0.3,0.5,0.7,0.9$, and three values of $r$ (cross symbols). In addition, the dependence of the analytical prediction for $t_{\rm min}$ from Eq. (35) on $\theta_{\mathrm{c}}$ and on $r$ is shown (solid lines). We observe that the analytical result yields a useful estimate for the dependence of the minimal validity time-scale on the model parameters and on the critical angle. We note that for very small values of $\theta_{\mathrm{c}}<0.1$, the analytical result for $t_{\rm min}(\theta_{\mathrm{c}})$ strongly increases. This is due to the fact that for such small values of the critical angle, only a very small range of scattering angles is treated via the drag force method, such that it formally leads to sufficiently frequent scatterings only after a rather long time. However, we notice that for $\theta_{\mathrm{c}}=0.1$ the value of $t_{\min}$ extracted from our numerical solutions is actually still rather small (left-most crosses in Fig. 7). We attribute this to the fact that the physical impact of scatterings with $\theta>\theta_{\mathrm{c}}$ is dominant for such small values of $\theta_{\mathrm{c}}$. Since those are treated based on the explicit sampling technique within hSIDM, the potentially larger relative error of the small-angle regime is not noticeable in the complete angular distribution function for $\theta_{\mathrm{c}}=0.1$.
@@
Thus, we find that the validity criteria for applying the effective small-angle approach derived in Sect. 3.2.3 are applicable in practice also beyond the strict set-up for which they have been derived, and yield a useful quantitative estimate for when the hSIDM approach can be used. Finally, we notice that $t_{\rm min}(\theta_{\mathrm{c}})$ is significantly smaller than the dynamical time-scale $t_{\rm dyn}(\equiv t_{1})$ as long as $\theta_{\mathrm{c}}$ is sufficiently small for both cases with (i) and without (ii) recoil. In addition, the ratio $t_{\rm min}/t_{\rm dyn}$ is the smaller the larger $r$. This implies the hSIDM approach is valid particularly for very strongly forward-dominated differential cross sections. Since the gain in numerical efficiency is also most pronounced for large $r$, this implies that hSIDM is indeed applicable in the cases for which it is expected to be most advantageous from the point of view of computational resources. We note that as an additional check, we explicitly verified that all results presented in Sect. 6 are insensitive to the choice of critical angle $\theta_{\mathrm{c}}$, being particularly relevant for moderate values of the anisotropy parameter $r\lesssim{\cal O}(10^{2})$.
@@ 5.4 Performance gain due to the hSIDM approach
Figure 8: Performance gain of the hybrid scheme (hSIDM) developed in this work for taking the angle-dependence of DM self-scattering (characteristic for e.g. light mediator models) into account in $N$-body simulations. We show the CPU run-time $t_{\rm cpu}$ (executed on a single CPU only) required for advancing the deflection test problem up to physical simulation time $t$ (in units of the typical reference time-scale $t_{1}$) in the upper panel. The red line corresponds to the naive scheme for which scatterings by all angles are sampled from $\mathrm{d}\sigma/\mathrm{d}\Omega$ (anisotropic rSIDM). The CPU-time within the hSIDM approach is shown for runs with various values of the critical angle $\theta_{\mathrm{c}}$ used to separate effective small- and explicit large-angle regimes. The CPU-time is reduced by almost a factor $100$ for $\theta_{\mathrm{c}}\sim 0.3$, and already significantly smaller for $\theta_{\mathrm{c}}\sim 0.1$. The lower panel shows the amount of CPU run-time $\Delta t_{\rm cpu}$ required to simulate a given physical simulation time interval $\Delta t_{\rm sim}$, which is roughly $t$-independent. The $y$-axis is normalised to unity for the anisotropic rSIDM scheme. The CPU run-time decreases strongly for $\theta_{\mathrm{c}}\lesssim 0.1$, and saturates for $\theta_{\mathrm{c}}\sim 0.3$. For even larger critical angles, the numerical cost of the effective small-angle approach becomes relevant as compared to the explicit treatment of large-angle scatterings within hSIDM.
@@
In summary, we verified the applicability and accuracy of the hSIDM approach, and demonstrated that it yields an improvement in computational performance by up to a factor $100$ for models of SIDM featuring a differential cross section with enhancement at small scattering angles, as is typical for light mediator models.
@@ 6 Simulation of merging galaxy clusters
Next, we describe the setup of the simulations, explain their analysis and demonstrate that the results are insensitive to the choice of the critical angle (Sect. 6.1). It follows a general description of the merger evolution (Sect. 6.2) as well as a discussion how well models of different angular dependencies can be mapped onto each other in the case of merging galaxy clusters (Sect. 6.3). A detailed investigation and discussion of the DM-galaxy offsets follow in Sect. 7.
@@ 6.1 Merger simulation set-up, analysis and validation
A characteristic signature of dark matter self-interactions are offsets between the dark matter and galaxy populations of the individual galaxy clusters, that are dynamically generated during the merger process. To quantify such offsets, it is necessary to identify a suitable ‘peak’ location of the dark matter and galaxy distributions, respectively, for which various methodologies have been explored (e.g., Power et al., 2003; Kim et al., 2017; Robertson et al., 2017b). Here we adopt the peak identification algorithm proposed by Fischer et al. (2021b), which tracks for each particle whether it belonged originally to the first or the second halo, [^15] We note that for Møller scattering the scattering angle $\theta$ larger than $\uppi/2$ are exchanged to $\uppi-\theta$ to ensure consistent association of the particles in the two halos. See also Sect. 2.1.2 by Fischer et al. (2021b). and utilises a peak search approach based on the gravitational potential. To compute the peak position for a component, for example, the galaxies of the first halo (association based on the ICs), the gravitational potential sourced by these particles only is computed and the peak position is set by the potential minimum. For the head-on collision set-up studied in this work, the peaks are always located along the merger axis. We denote the cluster moving initially in the negative (positive) direction of the merger axis as ‘halo 1’ (‘halo 2’). The peak positions along the merger axis are denoted by $d_{\rm DM}^{{\rm halo}\,n}$ and $d_{\rm Galaxies}^{{\rm halo}\,n}$ for the DM and galaxy populations and the two halos ($n=1,2$), respectively. In addition, we consider the location $d_{\rm BCG}^{{\rm halo}\,n}$ of the BCG for each cluster. We define the offsets as [^16] This definition matches the one used by (Fischer et al., 2021b), Fischer et al. (2023) and Sabarish et al. (2024) except that the offset has a different sign. We note that the definition by Fischer et al. (2021a) differs due to the use of absolute distances between the peaks of the same component.
@@
The hSIDM set-up involves the critical $\theta_{\mathrm{c}}$, that determines the separation between the regimes treated using the dedicated small-angle and large-angle approaches, respectively. As stated above, we checked that all our results are independent of the choice of $\theta_{\mathrm{c}}$ within numerical uncertainties. While we provide more details for specific quantities of interest below, we give a first example in Fig. 9, showing the one-dimensional density profile within a small region [^17] In practice we use a cuboid centred around the x-axis (merger axis) with depth of $100\,\mathrm{kpc}$ and width of $30\,\mathrm{kpc}$. around the merger axis, for various times, and for the DM density (solid lines) as well as the galaxy density (dashed lines). Here we used a strongly anisotropic Møller cross section with $r=10^{4}$. The results obtained from the hSIDM runs for $\theta_{\mathrm{c}}=0.1$ and $\theta_{\mathrm{c}}=0.3$ are shown by different colours, and are hardly distinguishable from each other in Fig. 9. This already indicates that peak positions and the resulting offsets between DM and galaxy distributions are independent of $\theta_{\mathrm{c}}$, as we will also verify later on.
We also stress that the hSIDM scheme allow us to simulate the strongly forward-dominated model for self-interactions described by $r=10^{4}$, which would be numerically prohibitively expensive when using an algorithm that does not employ an effective treatment of small-angle scatterings (i.e. for $\theta_{\mathrm{c}}=0$).
@@ 6.2 Peak positions
Figure 10: Time-evolution of the DM (red dashed) and galaxy (green dot-dashed) peak positions, as well as the BCG position (blue short dot-dashed), for both merging galaxy clusters. The upper left panel corresponds to CDM, while the other panels show three different models for SIDM: purely forward-dominated (fSIDM, upper right panel), isotropic (rSIDM, lower left panel), and a light mediator model described by a Møller cross section with $r=10$ (simulated using the hSIDM approach, lower right panel). The absolute self-interaction strength is matched in all cases by requiring a viscosity cross section $\sigma_{\mathrm{V}}/m_{\chi}=1\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$. In each panel we also show the peak positions shortly after the first pericentre passage (left inset) and those around apocentre (right inset).
In Fig. 10, we present the time-evolution of the peak positions of DM and galaxy distributions, as well as the BCG position, along the merger axis. The two sets of lines within each panel correspond to the first and second halo, respectively. For comparison, the upper left panel shows the purely collisionless DM case, that is without self-interactions. The peaks of DM and galaxies as well as the BCG position are then indistinguishable from each other for all times, as expected. Since we consider an equal-mass merger that occurs head-on, the peaks oscillate along the merger axis (around the centre of mass ) in a symmetrical way, with an amplitude that is damped due to dynamical friction. The other three panels in Fig. 10 show the evolution when assuming various models for DM self-interactions, specifically purely forward-dominated scattering (fSIDM), isotropic scattering (isotropic rSIDM), as well as an anisotropic Møller cross section with $r=10$ (hSIDM), respectively. We fixed the absolute self-interaction strength in all these cases such that they correspond to the same viscosity cross section, specifically $\sigma_{\mathrm{V}}/m_{\chi}=1\,{\rm cm}^{2}\,{\rm g}^{-1}$. While we discuss the rationale for this choice in more detail below (see Sect. 6.3), we already notice that it leads to an evolution of peak positions that is broadly similar in all cases. Nevertheless, significant differences between the models remain, see Sect. 7.2.
@@
6.3 Matching of models with different angular dependence?
@@
Figure 11: Comparison of the evolution of DM peak positions (left column), galaxy density peak positions (middle column) and the DM-galaxy offset (right column) for three different models considered in Sect. 6.3: almost isotropic model (1) in black, anisotropic model (2) with matched transfer cross section in orange and anisotropic model (3) with matched viscosity cross section in blue. In the lower panels the difference (2)-(1) (orange) and (3)-(1) (blue) is shown. While the evolution of the DM peak position can be approximately matched among (1) and (2), i.e. for constant transfer cross section, this is not the case for the evolution of galaxy positions and the DM-galaxy offset. In particular, the offset is larger for anisotropic models, by a factor $\sim 4$ ($\sim 2$) when comparing to an isotropic model with the same transfer (viscosity) cross section. Thus, a simulation including the angle-dependence is required in general to achieve reliable results.
The hSIDM approach allows us to simulate SIDM models with strongly anisotropic differential cross section. Using these results, we are able to scrutinise the question whether certain characteristic features of SIDM can be related among models with distinct differential cross sections by considering a suitable angle-averaged cross sections. Specifically, in this context, we are interested in the question whether the time-evolution of DM and galaxy peak positions during the merger process, as well as their offset, can be ‘mapped’ across different models. If such a mapping exists, it would imply that models with different differential cross section but identical angle-averaged cross section yield indistinguishable results for one particular quantity (or ideally even several) of interest.
@@
To be concrete, we consider three specific models, all of which we simulate with the hSIDM scheme:
@@
Let us now turn to the DM and galaxy peak positions, as well as their offset. The time-evolution is shown in the three columns of Fig. 11, respectively, with each panel including lines for model (1) in black, model (2) in orange, and model (3) in blue. In addition, the lower panels show the difference (2)-(1) (orange) and (3)-(1) (blue). From the left panel in Fig. 11, we see that the DM peak position for model (1) and (2) is similar at all times. This indicates that matching via the transfer cross section is a reasonable approximation for the evolution of the DM peaks. On the other hand, for the galaxy peak position (middle panel in Fig. 11) neither model (2) nor (3) are close to (1) at all times. However, during the most relevant stage from the first pericentre ($t\simeq 1.8$ Gyr) until around the first apocentre ($t\simeq 2.3$ Gyr), the galaxy peak position of model (3) is closer to model (1). This suggests that galaxy positions can be better matched using the viscosity cross section, but only until around the first apocentre. Afterwards, no clear matching seems possible.
@@
Finally, we turn to the offset between the DM and galaxy peak positions, shown in the right panel of Fig. 11. We observe that the offsets for both the anisotropic models (2) and (3) are significantly larger than for the isotropic model (1). In particular, this is the case shortly after the first pericentre passage at $t\simeq 2$ Gyr (when the DM peaks of the two clusters are closer to each other than the galaxy peaks, corresponding to negative offset) as well as around the first apocentre ($t\simeq 2.3$ Gyr, positive offset). The maximal absolute offsets in the two stages differ by a factor of around $2$ between model (3) and (1), and even by a factor of around $4$ for model (2) and (1). The finding of larger absolute values for the offsets in anisotropic models versus the isotropic case is in line with Fischer et al. (2021a). Thus we find that neither the transfer nor the viscosity cross section yield a useful matching between models when regarding the DM-galaxy offset during the most relevant stages of the merger evolution.
[…]
7 Model-dependence of DM-galaxy offsets

### v2

@@ 5.3 Validity check of hSIDM approach
Figure 6: Angular distribution $f(t,\theta)\cdot\sin(\theta)$ of beam particles within the deflection test set-up as obtained within the hSIDM scheme at three (very early) times $t$ (solid lines) and for two values of the critical angle, $\theta_{\mathrm{c}}=0.3$ (left panel) and $\theta_{\mathrm{c}}=0.5$ (right panel), respectively. The angular distribution is expected to agree with the analytical result Eq. (18) (dashed lines) after some minimal time $t>t_{\rm min}(\theta_{\mathrm{c}})$ for which scatterings by angles $\theta<\theta_{\mathrm{c}}$ are sufficiently frequent such that the effective drag force and momentum diffusion approach applies. The agreement is indeed reached after the expected time $t_{\rm min}(0.3)\simeq 2\cdot 10^{-2}t_{1}$ and $t_{\rm min}(0.5)\simeq 4\cdot 10^{-2}t_{1}$, respectively. We note that hSIDM is thus well applicable on the relevant dynamical timescale $t\sim t_{\rm dyn}(\equiv t_{1})$.
@@
Figure 7: Minimal timescale $t_{\rm min}(\theta_{\mathrm{c}})$ after which the hSIDM approach is applicable (extracted from the deflection test set-up). The symbols are derived from a comparison of numerical solutions using the hSIDM scheme with the result obtained when using the numerically expensive explicit sampling technique for all scattering angles (anisotropic rSIDM scheme). We show results for $\theta_{\mathrm{c}}=0.1,0.3,0.5,0.7,0.9$, and for differential cross sections with anisotropy parameter $r=10^{2},10^{3},10^{4}$, respectively. We furthermore consider the case without recoil and with ‘Rutherford’ scattering cross section (upper panel) as well we the case with recoil and with ‘Møller’ cross section (lower panel). Solid lines show the dependence of the analytical prediction Eq. (35) for $t_{\rm min}(\theta_{\mathrm{c}})$ on $\theta_{\mathrm{c}}$ and on $r$.
@@
We note that $t_{\rm min}(\theta_{\mathrm{c}})\ll t_{\rm dyn}(=t_{1})$ is much smaller than the dynamical timescale for both values of $\theta_{\mathrm{c}}$ shown in Fig. 6. This means the hSIDM approach yields a valid description on the relevant timescale of order $t_{\rm dyn}$. In the upper panel of Fig. 7 we show $t_{\rm min}(\theta_{\mathrm{c}})$ as obtained from hSIDM simulations with various critical angles $\theta_{\mathrm{c}}=0.1,0.3,0.5,0.7,0.9$, and three values of $r$ (cross symbols). In addition, the dependence of the analytical prediction for $t_{\rm min}$ from Eq. (35) on $\theta_{\mathrm{c}}$ and on $r$ is shown (solid lines). We observe that the analytical result yields a useful estimate for the dependence of the minimal validity timescale on the model parameters and on the critical angle. We note that for very small values of $\theta_{\mathrm{c}}<0.1$, the analytical result for $t_{\rm min}(\theta_{\mathrm{c}})$ strongly increases. This is due to the fact that for such small values of the critical angle, only a very small range of scattering angles is treated via the drag force method, such that it formally leads to sufficiently frequent scatterings only after a rather long time. However, we notice that for $\theta_{\mathrm{c}}=0.1$ the value of $t_{\min}$ extracted from our numerical solutions is actually still rather small (left-most crosses in Fig. 7). We attribute this to the fact that the physical impact of scatterings with $\theta>\theta_{\mathrm{c}}$ is dominant for such small values of $\theta_{\mathrm{c}}$. Since those are treated based on the explicit sampling technique within hSIDM, the potentially larger relative error of the small-angle regime is not noticeable in the complete angular distribution function for $\theta_{\mathrm{c}}=0.1$.
@@
Thus, we find that the validity criteria for applying the effective small-angle approach derived in Sect. 3.2.3 are applicable in practice also beyond the strict set-up for which they have been derived, and yield a useful quantitative estimate for when the hSIDM approach can be used. Finally, we notice that $t_{\rm min}(\theta_{\mathrm{c}})$ is significantly smaller than the dynamical timescale $t_{\rm dyn}(\equiv t_{1})$ as long as $\theta_{\mathrm{c}}$ is sufficiently small for both cases with (i) and without (ii) recoil. In addition, the ratio $t_{\rm min}/t_{\rm dyn}$ is the smaller the larger $r$. This implies the hSIDM approach is valid particularly for very strongly forward-dominated differential cross sections. Since the gain in numerical efficiency is also most pronounced for large $r$, this implies that hSIDM is indeed applicable in the cases for which it is expected to be most advantageous from the point of view of computational resources. We note that as an additional check, we explicitly verified that all results presented in Sect. 6 are insensitive to the choice of critical angle $\theta_{\mathrm{c}}$, being particularly relevant for moderate values of the anisotropy parameter $r\lesssim{\cal O}(10^{2})$.
@@ 5.4 Performance gain due to the hSIDM approach
Figure 8: Performance gain of the hybrid scheme (hSIDM) developed in this work for taking the angle dependence of DM self-scattering (characteristic for e.g. light mediator models) into account in $N$-body simulations. We show the CPU run-time $t_{\rm cpu}$ (executed on a single CPU only) required for advancing the deflection test problem up to physical simulation time $t$ (in units of the typical reference timescale $t_{1}$) in the upper panel. The red line corresponds to the naive scheme for which scatterings by all angles are sampled from $\mathrm{d}\sigma/\mathrm{d}\Omega$ (anisotropic rSIDM). The CPU-time within the hSIDM approach is shown for runs with various values of the critical angle $\theta_{\mathrm{c}}$ used to separate effective small- and explicit large-angle regimes. The CPU-time is reduced by almost a factor $100$ for $\theta_{\mathrm{c}}\sim 0.3$, and already significantly smaller for $\theta_{\mathrm{c}}\sim 0.1$. The lower panel shows the amount of CPU run-time $\Delta t_{\rm cpu}$ required to simulate a given physical simulation time interval $\Delta t_{\rm sim}$, which is roughly $t$-independent. The $y$-axis is normalised to unity for the anisotropic rSIDM scheme. The CPU run-time decreases strongly for $\theta_{\mathrm{c}}\lesssim 0.1$, and saturates for $\theta_{\mathrm{c}}\sim 0.3$. For even larger critical angles, the numerical cost of the effective small-angle approach becomes relevant as compared to the explicit treatment of large-angle scatterings within hSIDM.
@@
In summary, we verified the applicability and accuracy of the hSIDM approach. In particular, we demonstrated that it yields an improvement in computational performance by up to a factor $100$ for models of SIDM featuring a differential cross section with enhancement at small scattering angles, as is typical for light mediator models.
@@ 6 Simulation of merging galaxy clusters
Next, we describe the set-up of the simulations, explain their analysis and demonstrate that the results are insensitive to the choice of the critical angle (Sect. 6.1). It follows a general description of the merger evolution (Sect. 6.2) as well as a discussion how well models of different angular dependencies can be mapped onto each other in the case of merging galaxy clusters (Sect. 6.3). A detailed investigation and discussion of the DM-galaxy offsets follow in Sect. 7.
@@ 6.1 Merger simulation set-up, analysis and validation
A characteristic signature of dark matter self-interactions are offsets between the dark matter and galaxy populations of the individual galaxy clusters, that are dynamically generated during the merger process. To quantify such offsets, it is necessary to identify a suitable ‘peak’ location of the dark matter and galaxy distributions, respectively, for which various methodologies have been explored (e.g. Power et al., 2003; Kim et al., 2017; Robertson et al., 2017b). Here we adopt the peak identification algorithm proposed by Fischer et al. (2021b), which tracks for each particle whether it belonged originally to the first or the second halo, [^15] We note that for Møller scattering the scattering angle $\theta$ larger than $\uppi/2$ are exchanged to $\uppi-\theta$ to ensure consistent association of the particles in the two halos. See also Sect. 2.1.2 by Fischer et al. (2021b). and utilises a peak search approach based on the gravitational potential. To compute the peak position for a component, for example, the galaxies of the first halo (association based on the ICs), the gravitational potential sourced by these particles only is computed and the peak position is set by the potential minimum. For the head-on collision set-up studied in this work, the peaks are always located along the merger axis. We denote the cluster moving initially in the negative (positive) direction of the merger axis as ‘halo 1’ (‘halo 2’). The peak positions along the merger axis are denoted by $d_{\rm DM}^{{\rm halo}\,n}$ and $d_{\rm Galaxies}^{{\rm halo}\,n}$ for the DM and galaxy populations and the two halos ($n=1,2$), respectively. In addition, we consider the location $d_{\rm BCG}^{{\rm halo}\,n}$ of the BCG for each cluster. We define the offsets as [^16] This definition matches the one used by (Fischer et al., 2021b), Fischer et al. (2023) and Sabarish et al. (2024) except that the offset has a different sign. We note that the definition by Fischer et al. (2021a) differs due to the use of absolute distances between the peaks of the same component.
@@
The hSIDM set-up involves the critical $\theta_{\mathrm{c}}$, that determines the separation between the regimes treated using the dedicated small-angle and large-angle approaches, respectively. As stated above, we checked that all our results are independent of the choice of $\theta_{\mathrm{c}}$ within numerical uncertainties. While we provide more details for specific quantities of interest below, we give a first example in Fig. 9, showing the one-dimensional density profile within a small region [^17] In practice we use a cuboid centred around the x-axis (merger axis) with depth of $100\,\mathrm{kpc}$ and width of $30\,\mathrm{kpc}$. around the merger axis, for various times, and for the DM density (solid lines) as well as the galaxy density (dashed lines). Here we used a strongly anisotropic Møller cross section with $r=10^{4}$. The results obtained from the hSIDM runs for $\theta_{\mathrm{c}}=0.1$ and $\theta_{\mathrm{c}}=0.3$ are shown by different colours, and are hardly distinguishable from each other in Fig. 9. This already indicates that peak positions and the resulting offsets between DM and galaxy distributions are independent of $\theta_{\mathrm{c}}$, as we also verify later on.
We also stress that the hSIDM scheme allow us to simulate the strongly forward-dominated model for self-interactions described by $r=10^{4}$. It would be numerically prohibitively expensive when using an algorithm that does not employ an effective treatment of small-angle scatterings (i.e. for $\theta_{\mathrm{c}}=0$).
@@ 6.2 Peak positions
Figure 10: Time evolution of the DM (red dashed) and galaxy (green dot-dashed) peak positions, as well as the BCG position (blue short dot-dashed), for both merging galaxy clusters. The upper left panel corresponds to CDM, while the other panels show three different models for SIDM: purely forward-dominated (fSIDM, upper right panel), isotropic (rSIDM, lower left panel), and a light mediator model described by a Møller cross section with $r=10$ (simulated using the hSIDM approach, lower right panel). The absolute self-interaction strength is matched in all cases by requiring a viscosity cross section $\sigma_{\mathrm{V}}/m_{\chi}=1\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$. In each panel we also show the peak positions shortly after the first pericentre passage (left inset) and those around apocentre (right inset)
In Fig. 10 we present the time evolution of the peak positions of DM and galaxy distributions, as well as the BCG position, along the merger axis. The two sets of lines within each panel correspond to the first and second halo, respectively. For comparison, the upper left panel shows the purely collisionless DM case, that is, without self-interactions. The peaks of DM and galaxies as well as the BCG position are then indistinguishable from each other for all times, as expected. Since we consider an equal-mass merger that occurs head-on, the peaks oscillate along the merger axis (around the centre of mass ) in a symmetrical way, with an amplitude that is damped due to dynamical friction. The other three panels in Fig. 10 show the evolution when assuming various models for DM self-interactions, specifically purely forward-dominated scattering (fSIDM), isotropic scattering (isotropic rSIDM), as well as an anisotropic Møller cross section with $r=10$ (hSIDM), respectively. We fixed the absolute self-interaction strength in all these cases such that they correspond to the same viscosity cross section, specifically $\sigma_{\mathrm{V}}/m_{\chi}=1\,{\rm cm}^{2}\,{\rm g}^{-1}$. While we discuss the rationale for this choice in more detail below (see Sect. 6.3), we already notice that it leads to an evolution of peak positions that is broadly similar in all cases. Nevertheless, significant differences between the models remain, see Sect. 7.2.
@@
6.3 Matching of models with different angular dependence
@@
Figure 11: Comparison of the evolution of DM peak positions (left column), galaxy density peak positions (middle column), and the DM-galaxy offset (right column) for three different models considered in Sect. 6.3. The result for the almost isotropic model (1) is shown in black, the result for the anisotropic model (2) with matched transfer cross section is shown in orange, and the result for the anisotropic model (3) with matched viscosity cross section is shown in blue. In the lower panels the difference (2)-(1) (orange) and (3)-(1) (blue) is shown. While the evolution of the DM peak position can be approximately matched among (1) and (2), i.e. for constant transfer cross section, this is not the case for the evolution of galaxy positions and the DM-galaxy offset. In particular, the offset is larger for anisotropic models, by a factor of $\sim 4$ ($\sim 2$) when comparing to an isotropic model with the same transfer (viscosity) cross section. Thus, a simulation including the angle dependence is required in general to achieve reliable results.
The hSIDM approach allows us to simulate SIDM models with strongly anisotropic differential cross section. Using these results, we are able to scrutinise the question whether certain characteristic features of SIDM can be related among models with distinct differential cross sections by considering a suitable angle-averaged cross sections. Specifically, in this context, we are interested in the question whether the time evolution of DM and galaxy peak positions during the merger process, as well as their offset, can be ‘mapped’ across different models. If such a mapping exists, it would imply that models with different differential cross section but identical angle-averaged cross section yield indistinguishable results for one particular quantity (or ideally even several) of interest.
@@
To be concrete, we considered three specific models, and we simulated all of them with the hSIDM scheme:
@@
Let us now turn to the DM and galaxy peak positions, as well as their offset. The time evolution is shown in the three columns of Fig. 11, respectively, with each panel including lines for model (1) in black, model (2) in orange, and model (3) in blue. In addition, the lower panels show the difference (2)-(1) (orange) and (3)-(1) (blue). From the left panel in Fig. 11, we see that the DM peak position for model (1) and (2) is similar at all times. This indicates that matching via the transfer cross section is a reasonable approximation for the evolution of the DM peaks. On the other hand, for the galaxy peak position (middle panel in Fig. 11) neither model (2) nor (3) are close to (1) at all times. However, during the most relevant stage from the first pericentre ($t\simeq 1.8$ Gyr) until around the first apocentre ($t\simeq 2.3$ Gyr), the galaxy peak position of model (3) is closer to model (1). This suggests that galaxy positions can be better matched using the viscosity cross section, but only until around the first apocentre. Afterwards, no clear matching seems possible.
@@
Finally, we turn to the offset between the DM and galaxy peak positions, shown in the right panel of Fig. 11. We observe that the offsets for both the anisotropic models (2) and (3) are significantly larger than for the isotropic model (1). In particular, this is the case shortly after the first pericentre passage at $t\simeq 2$ Gyr (when the DM peaks of the two clusters are closer to each other than the galaxy peaks, corresponding to negative offset) as well as around the first apocentre ($t\simeq 2.3$ Gyr, positive offset). The maximal absolute offsets in the two stages differ by a factor of around two between model (3) and (1), and even by a factor of around four for model (2) and (1). The finding of larger absolute values for the offsets in anisotropic models versus the isotropic case is in line with Fischer et al. (2021a). Thus we find that neither the transfer nor the viscosity cross section yield a useful matching between models when regarding the DM-galaxy offset during the most relevant stages of the merger evolution.
In conclusion, while an approximate mapping of DM peak positions with the transfer cross section is reasonable, the galaxy positions as well as the DM-galaxy offset cannot be mapped with this prescription, with offsets differing by a factor of four between isotropic and strongly anisotropic cases. When comparing models with different angular dependence but identical viscosity cross section, we observe that the galaxy peak position can be roughly mapped until about the first apocentre. However, in that case the DM peak positions differ for the isotropic and anisotropic models, such that the DM-galaxy offsets differ by a factor of around two. Altogether, this implies that the complex merger dynamics requires to include the actual angle dependence for a given model of self-interactions in order to achieve reliable predictions.
7 Model dependence of dark matter-galaxy offsets